Production costs chargeable to the Finishing Department in June in Hollins Company are materials $16,880, labor $24,006, overhead $18,500. Equivalent units of production are materials 21,100 and conversion costs 20,050. Production records indicate that 19,000 units were transferred out, and 2,100 units in ending work in process were 50% complete as to conversion costs and 100% complete as to materials.

Prepare a cost reconciliation schedule.
Answer #1
Cost Reconciliation
Cost accounted for:
Transferred out 55480 =19000*(0.80+2.12)
Work In process
Materials 1680 =2100*0.80
Conversion costs 2226 3906 =2100*50%*2.12
Total Cost accounted for 59386
Workings:
Materials Conversion costs
Total costs 16880 42506
Divide by Equivalent units 21100 20050
Cost per Equivalent unit 0.80 2.12